OpenStudy (anonymous):

Help: Which equation represents this line in point-slope form? (y−6)=−7/9(x+4) (y−3)=7/9(x−6) (y−3)=−9/7(x−3) (y−3)=9/7(x−3)

OpenStudy (anonymous):

which one do you think?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

look, to find the correct point slope form of this this line, first find the slope, of the line. to find the slope, use the slope formula \[m=\frac{ y_2-y_2 }{ x_2-x_1 }\] the slope of your line will be 9/7. So already there is only d is correct. But if you want to know how to continue the finding of point slope form of this line, here it is: (could you just wait a minute, please? it will take some time to me to write the explanation)

OpenStudy (anonymous):

the point slope form of the line is \[y-y_1=m(x-x_1)\] or \[y-y_2=m(x-x_2)\]

OpenStudy (anonymous):

since the line goes through points (3, 3) and (-4, -6), the point slope form of the line will be the following: \[y-3=9/7(x-3)\] AND \[y-(-6)=9/7(x-(-4))\] \[y+6=9/7(x+4)\] Note that both equation are correct, but in this problem you have only one equation. which is the last one. (in your choices)

OpenStudy (anonymous):

oh ok I got everything you said thank you so much!! @ksanka

OpenStudy (anonymous):

np :)